Enterprise, Inc. bonds have an annual coupon rate of 16 percent. The interest is paid semiannually and the bonds mature in 15 years. Their par value is ​$1 comma 000. If the​ market's required yield to maturity on a​ comparable-risk bond is 11 ​percent, what is the value of the​ bond? What is its value if the interest is paid​ annually? a. The value of the Enterprise bonds if the interest is paid semiannually is ​$ nothing. ​(Round to the nearest​ cent.)
Answer #1

Solution:

Coupon rate = 16%

PMT = 1000*16%/2= $80

Years to maturity = 15 years

Nper = 15*2= 30

Par value ( FV) = $1000

Yield to maturity = 11%

Rate = 11%/2= 5.5%

Present value ( PV) = ?

If interest is paid semiannually :-

PV ( rate, nper, PMT, FV)

PV ( 5.5%, 30, 80, 1000)

PV = $1363.34

If interest is paid annually

PV ( 11%, 15,160,1000)

PV = $1359.54